### v1

# One Model to Train them All: Hierarchical Self-Distillation for Enhanced Early Layer Embeddings

Abstract
Deploying language models often requires handling model size vs. performance trade-offs to satisfy downstream latency constraints while preserving the model’s usefulness. Model distillation is commonly employed to reduce model size while maintaining acceptable performance. However, distillation can be inefficient since it involves multiple training steps. In this work, we introduce ModularStarEncoder, a modular multi-exit encoder with 1B parameters, useful for multiple tasks within the scope of code retrieval. ModularStarEncoder is trained with a novel self-distillation mechanism that significantly improves lower-layer representations—allowing different portions of the model to be used while still maintaining a good trade-off in terms of performance. Our architecture focuses on enhancing text-to-code and code-to-code search by systematically capturing syntactic and semantic structures across multiple levels of representation. Specific encoder layers are targeted as exit heads, allowing higher layers to guide earlier layers during training. This self-distillation effect improves intermediate representations, increasing retrieval recall at no extra training cost. In addition to the multi-exit scheme, our approach integrates a repository-level contextual loss that maximally utilizes the training context window, further enhancing the learned representations. We also release a new dataset constructed via code translation, seamlessly expanding traditional text-to-code benchmarks with code-to-code pairs across diverse programming languages. Experimental results highlight the benefits of self-distillation through multi-exit supervision.
One Model to Train them All: Hierarchical Self-Distillation for Enhanced
Early Layer Embeddings
Andrea Gurioli1, Federico Pennino1, Joao Monteiro2 Maurizio Gabbrielli1 1University of Bologna, 2Autodesk Correspondence: andrea.gurioli5@unibo.it, joao.monteiro@autodesk.com
1 Introduction
Large language models (LLMs) have significantly impacted the field of natural language processing, demonstrating remarkable performance across various applications (Niu et al., 2023). However, the amount of computation required to operate state-of-the-art models poses significant challenges for the large-scale deployment of these models.
To mitigate these challenges, the research community has explored several model strategies to reduce the operational cost of LLMs without sacrificing their effectiveness. A prominent technique in model compression is quantization (Jacob et al., 2017; Lin et al., 2023; Egiazarian et al., 2024), which involves the reduction of numerical precision in the model’s parameters. Quantization effectively decreases memory requirements and enhances inference speed, facilitating the deployment of large language models in resource-constrained environments. Concurrently, knowledge distillation has emerged as a powerful technique whereby a smaller “student” model is trained to emulate the behavior of a larger “teacher” model, as evidenced by works such as DistilBERT (Sanh et al., 2019) and TinyBERT (Jiao et al., 2019). Additionally, pruning methods selectively eliminate less influential weights or neurons, further reducing model complexity and aiming to preserve performance (Han et al., 2015).
Recent efforts have increasingly focused on developing efficient architectures requiring fewer parameters. Model families such as LLaMA (Dubey et al., 2024), Qwen (Hui et al., 2024), Mistral (Jiang et al., 2023), and SmolLM (Allal et al., 2025) exemplify a paradigm shift towards smaller, more accessible architectures. These model families are deployed at various resolutions—ranging from lightweight variants optimized for heavily resource-constrained environments to larger versions that retain competitive performance.
Figure 1: Overview of our multi-exit self-distillation encoder, shown here with exit heads at selected layers (e.g., Layers 4, 9, 18, 27, and 36). Each exit head predicts an output embedding and adds a “layer loss,” contribution weighted by a coefficient $\alpha_{i}$, summed into the overall objective $\mathcal{L}$.
In parallel, advancements in dynamic inference strategies have introduced mechanisms that further optimize computational efficiency. Techniques like multi-exit networks enable early predictions at intermediate layers, reducing unnecessary computations. For instance, early-exit architectures such as BranchyNet (Teerapittayanon et al., 2017) dynamically balance computation and accuracy by allowing predictions before full model execution. Similarly, Matryoshka representation learning (Kusupati et al., 2022) extends this idea to embeddings, introducing a loss function that yields multi-granular representations. This approach allows downstream tasks to adjust computational complexity by pruning embedding dimensionality, further contributing to efficient model deployment.
Building on these principles, we propose ModularStarEncoder, a modular multi-exit encoder architecture that integrates a novel intra-model self-distillation mechanism. In our design, specific intermediate layers are supervised by both the primary task loss and auxiliary distillation losses on specific exit heads, encouraging lower layers to learn better representations by mimicking the outputs of higher layers. We apply a shared embedding head comprising a masked language modeling head and an in-context classification head across a chosen subset of layers. We then fine-tuned the model with different projection heads for each exit point. We reached state-of-the-art results on multiple retrieval tasks (such as code-to-code and text-to-code), fine-tuning one single modular model that can be sliced depending on the end-user computational constraints.
Our contributions are as follows:
  • We introduce a self-distillation framework that enables training multiple model resolutions within a unified layer stack, reducing redundancy and improving scalability. We believe this approach can significantly affect LLM training pipelines that depend on multiple model distillations.
  • We train and release ModularStarEncoder, which consists of a pre-trained [^1] https://huggingface.co/modularStarEncoder/ModularStarEncoder and fine-tuned [^2] https://huggingface.co/modularStarEncoder/ModularStarEncoder-finetuned encoder: The former is a modular pre-trained encoder with up to 1 billion parameters and five exit points, allowing users to perform multiple exit fine-tuning depending on downstream tasks. The latter is a fine-tuned encoder for various retrieval tasks. We allow the user to choose either the entire model with 1 billion parameters or a model size that fits their memory and computational limitations.
  • We release SynthCode2Code2NL a new dataset [^3] https://huggingface.co/datasets/modularStarEncoder/SynthCode2Code2NL constructed via code translation, expanding popular text-to-code datasets across diverse programming languages with code-to-code pairs. SynthCode2Code2NL comprises 1 071 367 triplets of natural language-code-code.
2 Methodology
2.1 Dataset
Table 1: SynthCode2Code2NL details: Average character count and sample size per language for the CodeSearchNet dataset and the synthesized portion obtained through translation.
Language CSN samples CSN avg. char Synth. samples Synth. avg. char
English 1 071 367 180 - -
PHP 280 706 514 116 967 579
Python 274 454 474 117 374 518
Go 234 089 350 124 125 541
Java 282 118 505 116 098 707
C++ - - 141 956 938
Ruby - - 158 494 456
C - - 136 365 1029
JavaScript - - 159 988 557
Figure 2: Prompt provided to Qwen2.5-Coder-7B-Instruct for translating a given code snippet ( print("Hello World") in the example) from a source programming language (Python) to a target one (Rust).
Table 2: Hyperparameters for Architecture, Pre-training, and Fine-tuning
Architecture
Hyperparameter Value
Model size 1B parameters
Precision bfloat16
Hidden layers 36
Attention heads 16
Hidden dimensionality 1024
Positional encoding RoPE ($\theta=10^{-6}$)
Context length 2048
Attention mechanism Grouped-Query Attention
Attention pattern Bi-directional
Pre-training
Batch size 3.99M tokens
Pretraining steps 245 000
Pretraining Tokens 1T
Loss function MLM + In-Context loss
Multi-layer loss yes
Optimizer AdamW
Weight decay 1e-1
Initial learning rate 6.24e-4
Learning rate schedule Multi-step
Warmup steps 4000
Fine-tuning
Dataset size 635.404 samples
Fine-tuning steps 20 000
Loss function CLIP loss
Multi-layer loss yes
Batch size 2048
Learning rate 1.0e-5
Temperature parameter 10.0
Hardware (Pre-training + fine-tuning)
GPUs 512 NVIDIA Ampere (64GB)
Overall Training hours 450 000
In the pre-training phase, we leveraged The Stack V2 Lozhkov et al. (2024), a large open-source code dataset structured by repository.
For the fine-tuning stage, we created SynthCode2Code2NL, a dataset that supports text-to-code and code-to-code search. Using the popular CodeSearchNet Husain et al. (2019) as a seed dataset and selecting popular programming languages (Python, Java, Go, and PHP), we augmented it by transpiling available code snippets onto other languages.
To generate semantically similar code snippets for code-to-code search, we translated each snippet into a different language randomly sampled from Go, Ruby, Python, Java, C++, PHP, C, JavaScript. We prompted the Qwen2.5-Coder-7B-Instruct model with the source code, the name of the source language, and the name of the target language (see fig. 2). During code translation, we choose the token with the highest probability as output (greedy search) to prevent semantic discrepancies.
This process yielded pairs of code snippets in distinct languages tied to the same natural language description. As a result, every sample in the fine-tuning dataset includes a natural language description and two code snippets from distinct languages. SynthCode2Code2NL contains 1 071 367 samples where, in the first code column, we directly processed code snippets from CodeSearchNet, including Python, Java, PHP, and Go. The third column, artificially synthesized via code translation, includes Go, Ruby, JavaScript, Python, C++, PHP, C, and Java code snippets. After a manual inspection, we discovered that both columns contained code snippets that differed only in identifiers or function arguments. Several tasks were semantically identical but paraphrased with different parameter requirements (e.g., two identical paraphrased tasks asked for opening a socket on a different port). During the preprocessing phase of SynthCode2Code2NL, motivated by the dataset’s redundancy and preliminary experiments that show its effectiveness on the model’s performance, we near-deduplicated the dataset using both the CodeSearchNet code column and the synthesized code column. During the data near deduplication phase, we relied on Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 and 256 permutations, analyzing character-level 5-grams. Table 1 shows the average number of characters per language in SynthCode2Code2NL, we emphasize that synthesized data is significantly longer than human-written code and might have stylistic differences compared to human code, we further discussed this in Limitations. Appendix A provides examples of code translation.
Figure 3: On the left side the illustration of the in-context loss framework, where samples from different repositories are concatenated. Positive examples share the same repository context, while negative examples come from different repositories. On the right side, in-context loss framework pseudocode.
2.2 Architecture
We updated the first version of StarEncoder (Li et al., 2023) by enabling longer code snippets (up to 2 048 tokens as context length), increasing the model size from $\approx$125M to $\approx$1B parameters and utilizing state-of-the-art methodologies (Warner et al., 2024; Lozhkov et al., 2024) resulting in ModularStarEncoder.
We built ModularStarEncoder on top of StarCoder-2 (Lozhkov et al., 2024), applying several modifications to the model. We reduced its size from 15B to 1B parameters. Our architecture comprises 36 hidden layers and adopts Grouped Query Attention (GQA) (Ainslie et al., 2023) with 16 attention heads and 4 key-value heads. ModularStarEncoder relies upon Rotary Positional Encoding (RoPE) (Su et al., 2021) with a base period $\theta=10^{-6}$ and features a hidden dimensionality of 1024 with an intermediate size of 12 288.
We followed Devlin et al. (2019) and replaced the causal self-attention in StarCoder-2 with bidirectional self-attention. Aiming for modularity, we also replaced sliding window attention with full attention. This step was taken to avoid the receptive field phenomenon of sliding window mechanisms (Zhu et al., 2021). Finally, our implementation integrates FlashAttention V2 (Dao, 2023) for faster inference. Table 2 summarizes the architectural details.
2.3 Pre-training
We pre-trained ModularStarEncoder with a batch size of 3.99M tokens for 245 000 training steps, processing $\approx$1T tokens. We conducted pre-training and fine-tuning on 512 NVIDIA Ampere (64GB) GPUs using the Leonardo supercomputer (Turisini et al., 2023), requiring 450 000 GPU working hours.
To enable both token-level and snippet-level embeddings after pre-training, we employed a multi-objective pre-training strategy that combined two losses, as detailed in section 2.3.1 and section 2.3.2. The pre-training was performed on TheStackV2, whose context length analysis revealed an average of $\approx 630$ tokens per code snippet. As described in section 2.3.1, we concatenated multiple snippets to facilitate our multi-loss methodology, allowing our in-context classification loss to expand the average context window to $\approx 1300$ tokens, reaching the maximum context length 20% of the time.
We used the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ to 0.95, $\epsilon$ to 1e-6, and a weight decay of 1e-1. We initialized the learning rate at 6.24e-4 and decreased it using a multi-step learning rate schduler Bi et al. (2024) with 4 000 warmup steps. The learning rate was reduced at 120 000, 185 000, 220 000, 230 000, and 240 000 training steps, applying a decay factor of 0.36, and from step 185,000 onward, further reduced by factors of 0.1, 0.031, 0.01, and 0.001. Table 2 summarizes the hyperparameters for architecture, pre-training, and fine-tuning.
2.3.1 Masked Language Modeling and In-Context Classification
The training objectives of BERT (Feng et al., 2020), specifically Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), have become a de facto standard. However, The NSP loss constrains the context window length to the sentence length, leading to too many padding tokens and redundant computation (Zeng et al., 2022), and has been shown to not yield significant benefits after fine-tuning (Warner et al., 2024; Aroca-Ouellette and Rudzicz, 2020). Given that the average number of tokens per data sample in Stack v2 is 630, a large context window of 2048 results in substantial padding, making long-context training inefficient. While Wang et al. (2023) demonstrated the advantages of training LLMs with multiple objectives, we revisited the NSP loss and introduced an in-context classification (ICC) objective. We hypothesize that predicting whether multiple code snippets belong to the same context (in our case, the same repository) can enhance semantic search performance while allowing efficient concatenation of multiple code fragments. Our final training objective is the summation of two losses: (1) MLM loss and (2) ICC loss: $\mathcal{L}=\mathcal{L}_{MLM}+\mathcal{L}_{ICC}$.
In $\mathcal{L}_{MLM}$, a certain percentage of tokens are randomly masked and predicted using a classification head. Following Zhang et al. (2024), we adopt a 15% masking rate with the standard 80-10-10 token replacement strategy Devlin et al. (2019). The secondary objective, $\mathcal{L}_{ICC}$, determines whether randomly concatenated inputs (separated by a $<SEP>$ token) originate from the same repository (see fig. 3). Each concatenated sample has a 50% probability of containing source code from different repositories. This approach increases input density—reducing padding by expanding the average input length from 630 to 1 300 tokens—and potentially enhances cross-language understanding. Since repositories are inherently modular and often contain files written in multiple languages, learning from repository-level context may improve inter-language generalization.
Table 3: Performance of different models on text-to-code with CodeSearchNet using codeXGLUE. We reported the results presented in codet5plus, unixcoder and modernBERT (Wang et al., 2023; Guo et al., 2022; Warner et al., 2024).
CodeSearchNet
Model Ruby JS Go Python Java PHP avg. MRR avg. NDCG
ModularStarEncoder 74.1 74.0 82.5 92.5 78.7 84.5 81.0 84.2
Codet5+ 770M 78.0 71.3 92.7 75.8 76.2 70.1 77.4 -
OpenAI text-embedding-3-large 84.7 85.3 95.9 99.8 90.1 95.6 91.9 93.3
Unixcoder 74.0 68.4 91.5 72.0 72.6 67.6 74.4 -
ModernBERT-large - - - - - - - 59.5
2.3.2 Multi-layer Loss
To achieve layer-wise modularity in transformer architectures, we apply the previously introduced loss (section 2.3.1) across a selected set of layers, sharing classification heads (masked language modeling and in-context classification) while incorporating a positional embedding of the layer index. The total loss is computed as the sum of individual layer losses, weighted by a factor $\alpha$ to prioritize deeper layers: $\mathcal{L}=\sum_{i\in\iota}\mathcal{L}_{i}\cdot\alpha$ where $\alpha=i/|I|$ and $I=\{1,\dots,36\}$ represents all layers, and the selected subset $\iota=\{4,9,18,27,36\}$ defines the layers where the loss is applied. The selected subset was chosen to enable four model variants equally spaced in depth (9, 18, 27, 36) along with an additional “tiny” version (4) to see the model performance in a lower number of parameters set. This approach allows for flexible model deployment, enabling adaptive layer pruning while maintaining performance trade-offs.
2.4 Fine-tuning
Following Su et al. (2023), we fine-tune a single model for both text-to-code and code-to-code retrieval using instruction prompting. The optimization objective combines CLIP loss (Radford et al., 2021) with a multi-layer loss (details in 2.3.2).
To enhance representation learning, we replace the single-head projection of the multi-layer loss with five distinct projection heads, applied at different exit points of the pre-trained model (layers 4, 9, 18, 27, and 36). We used a batch of 2 048 elements, ensuring that text-to-code and code-to-code were equally distributed across the batch.
We performed data augmentation by randomly replacing frequently occurring words (appearing more than twice and having at least three characters) with random strings. We applied the augmentation exclusively to code snippets in 30% of cases, leaving natural language descriptions unchanged. After conducting a grid search, we selected $51.0e-$ 5 as the learning rate, maintained throughout the finetuning process, and set the temperature parameter at 10.0.
2.5 Evaluation
We evaluated ModularStarEncoder fine-tuned, on both text-to-code and code-to-code retrieval tasks using CodeXGLUE (Lu et al., 2021), which comprises several benchmarking datasets. For text-to-code retrieval, we employed the CodeSearchNet dataset, where the goal is to retrieve the most relevant code snippet given a natural language query. Specifically, the query corresponds to a documentation comment, and the model is tasked with ranking the correct code snippet among 999 distractor snippets Husain et al. (2019). This setup assesses the model’s ability to learn meaningful cross-modal representations between code and natural language.
For code-to-code retrieval, we relied on two datasets from CodeXGLUE: the Code Translation (CT) benchmark and POJ-104. The Code Translation dataset consists of semantically equivalent code snippets in different programming languages, and we framed the task as cross-language code retrieval rather than translation. In this setting, given a Java code snippet as a query, the model retrieves the corresponding C# implementation, testing its capability to capture cross-lingual semantic similarities between functionally equivalent programs.
In contrast, withPOJ-104 dataset, we want to evaluate the model on intra-language semantic search (POJ-104 contains only C++ snippets), where programs solve the same problem but with different implementations. This setup evaluates the model’s capacity to generalize across structural variations while preserving semantic equivalence.
Table 4: Performance of different models on Code Translation (CT) and POJ104 for code-to-code search with codeXGLUE dataset.
CT POJ104
Model MRR mAP
ModularStarEncoder 98.9 56.5
Codet5+ 110M-embedding 98.4 24.5
OpenAI text-embedding-3-large 98.8 82.9
Unixcoder 97.6 41.0
ModernBERT-large 93.1 27.3
3 Results and Discussion
Table 5: Performance comparison of ModularStarEncoder layers and baseline fine-tuned models on the CodeSearchNet benchmark. The table displays the overall retrieval performance measured by Mean Reciprocal Rank (MRR). We refer to ModularStarEncoder, fine-tuned with multiple exit points simultaneously, as self-distilled. The models not marked as self-distilled are the baselines, fine-tuned individually for each exit point.
CodeSearchNet
Model Size Ruby Javascript Go Python Java PHP avg. MRR
Layer-4 $\approx 160M$ 59.5 61.3 72.1 86.2 68.2 75.5 70.5
Layer-4 (self-distilled) 62.2 64.7 74.8 88.1 71.4 78.0 73.2
Layer-9 $\approx 300M$ 64.9 65.7 74.3 87.3 72.0 78.8 73.8
Layer-9 (self-distilled) 67.6 69.4 78.9 90.2 75.5 82.3 77.3
Layer-18 $\approx 550M$ 73.8 73.5 82.4 92.1 78.4 84.0 80.7
Layer-18 (self-distilled) 74.1 74.0 82.5 92.5 78.7 84.5 81.0
Layer-27 $\approx 800M$ 72.3 71.8 80.8 90.8 76.9 82.3 79.1
Layer-27 (self-distilled) 73.2 73.3 81.7 92.1 77.8 83.8 80.3
Layer-36 $\approx 1B$ 72.3 72.9 80.7 91.5 77.1 82.9 79.5
Layer-36 (self-distilled) 73.5 72.6 80.5 91.4 76.9 82.7 79.6
3.1 Benchmarks
Table 3 presents the results for CodeSearchNet (t2c) task in terms of Mean Reciprocal Rank (MRR) for each single language, average NDCG and average MRR. Results for Unixcoder, ModernBERT, and CodeT5+ are reported from the original papers (Guo et al., 2022; Warner et al., 2024; Wang et al., 2023). On CodeSearchNet, ModularStarEncoder achieves an MRR of 81.0 and a NDCG of 84.2, outperforming CodeT5+ (Wang et al., 2023) (770M), UniXcoder (Guo et al., 2022), and ModernBERT-large (Warner et al., 2024). The only encoder that surpasses ModularStarEncoder is OpenAI’s text-embedding-3-large.
Table 4 presents results from both POJ104 and CT datasets reported respectively in MRR for code translation (Java to C# retrieval) and mean average precision for POJ104 (C++ to C++# retrieval). ModularStarEncoder reaches the best performance among the tests. We decided to replicate the benchmarking for all models in a zero-shot setting for code-to-code tasks because our model does not integrate POJ104 and the code translation datasets in the training set.
Referring to Table 4, on the POJ104 dataset in zero-shot, ModularStarEncoder achieves an mAP of 0.57, which is state-of-art between open-sourced models, however it is significantly behind OpenAI text-embedding-3-large. We underscore that a direct comparison with OpenAI text-embedding-3-large remains challenging because it is closed-source, and details such as model size, training methodology, or potential data contamination are undisclosed.
3.2 Ablation Study
We conducted an ablation study by fine-tuning singularly each exit point (also starting from ModularStarEncoder, pre-trained) and pruning the subsequent layers (e.g., for the baseline on layer 18, we retain only the first 18 layers and fine-tune the model using just one projection head on that layer). Finally, we compared the sliced models with the corresponding results (self-distilled) of the model fine-tuned with the multi-layer loss (ModularStarEncoder). ModularStarEncoder consistently outperforms the single-exit baseline, indicating that lower-level layers benefit from training signals propagated from deeper layers. This behavior is highlighted in Table 5, where ModularStarEncoder, indicated as self-distilled, outperforms all the single exit baselines consistently. This finding underscores a promising new direction in self-distillation for large-scale code and text models, enabling high performance even in more compact configurations. Moreover, Figure 4 illustrates that ModularStarEncoder maintains robust performance from layers 18 to 36, allowing users to scale down the network to match their memory, computational, or latency constraints while preserving strong retrieval accuracy.
4 Related work
Since the introduction of ELMo (Peters et al., 2018), deep contextual information has enhanced generating embeddings for textual retrieval or classification, reaching state-of-the-art results in several tasks. BERT (Feng et al., 2020) followed those findings, adapting the Transformer architecture (Vaswani et al., 2017) to enable a bi-directional representation with two different training objecting, namely the masked language modeling and the next sentence prediction losses. Lan et al. (2019); Liu et al. (2019) adapted the BERT architecture to obtain an enhanced pre-trained model by removing or modifying the NSP, focusing on pre-training data or hyperparameters optimization. More recently, modernBERT (Warner et al., 2024) tied the gap between modern decoders (Jiang et al., 2023; Hui et al., 2024; Dubey et al., 2024; Touvron et al., 2023; Lozhkov et al., 2024) advancements that rely upon models with an increased number of parameters, trained upon more tokens, and being capable of handling longer contextual information.
(a) MRR
(b) Recall@1
Figure 4: Performance Comparison Across Layers: The graph illustrates the MRR and the Recall@1 for different layers, comparing baseline models and a self-distilled model.
In code representation, large language models must be adapted by training them on a curated corpus focused on software and by leveraging code’s syntactic and semantic structures, which differ significantly from natural language. Feng et al. (2020) adapted the BERT architecture to produce semantically meaningful embeddings for source code, resulting in codeBERT. This was accomplished by including more source code in the training set and focusing on a training loss that can leverage bimodal (natural language and code) contextual information (Clark et al., 2020). GraphCodeBERT enhanced codeBERT (Feng et al., 2020) representations by incorporating data flow graphs, capturing dependencies between variables and operations, and improving tasks like code summarization and clone detection. UniXcoder (Guo et al., 2022) extended this by introducing a unified encoder-decoder framework, integrating abstract syntax trees (ASTs) and data flow information. Wang et al. (2023) expanded these findings with codet5plus, stressing how multiple losses that leverage code semantics impact the model pertaining. The work incorporated text-code contrastive learning, text-code Matching, and text-code causal LM for better code understanding and generation.
When trying to achieve better performance, research has shifted toward models with a high number of parameters. While this trend appears effective from a performance perspective, end users may face computational or memory limitations as LLMs vary from millions to billions of parameters. Sanh et al. (2019) pioneered the introduction of knowledge distillation, using a “teacher” model that guides a smaller model to emulate its behavior. This methodology has been widely adopted and improved upon recently (DeepSeek-AI et al., 2025; Hui et al., 2024), becoming a standard for obtaining high-performing smaller LLMs.
Our work differs from previous work by adapting a modern architecture Lozhkov et al. (2024) to a code encoder-only based model and introducing a novel ’self-distillation’ mechanism. We replace the next sentence prediction loss with an in-context classification focused on the repository level and expand the context to 2048 tokens. Our novel self-distillation mechanism improves low-level layers, resulting in a modular transformer architecture without additional teacher models or further data for distillation.
5 Conclusion
In this work, we introduced ModularStarEncoder, a modular multi-exit encoder architecture designed to improve efficiency and scalability in code retrieval tasks. By integrating an intra-model self-distillation mechanism, our approach enables multiple resolution models to be trained within a unified layer stack, reducing redundancy while maintaining high retrieval performance. Our evaluation on CodeXGLUE demonstrates that ModularStarEncoder achieves state-of-the-art results among open-source models, outperforming prior baselines across text-to-code and code-to-code retrieval tasks. Ablations further highlighted the benefits of self-distillation, showing that lower layers gain representational strength from deeper layers, leading to superior performance compared to single-exit models.
Beyond performance gains, ModularStarEncoder offers practical benefits by providing multiple exit points, allowing users to balance computational efficiency and accuracy based on resource constraints. The results suggest that self-distillation provides a promising direction for efficient large-scale encoders, reducing deployment costs without sacrificing effectiveness.
Finally, released in open-access our SynthCode2Code2NL and both pre-trained and fine-tuned ModularStarEncoder models.
Acknowledgments
We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).
Limitations
Due to our dependence on multiple GPUs, we encountered significant computational constraints. Parameter grid searches with smaller and embryonic models were the only ways to extrapolate the best hyperparameter setup. The best hyperparameters for smaller models can differ from those for larger ones; thus, we faced a limitation in finding an optimal training setup. Ablating both the in-context classification and the multi-layer loss in a real scenario was impossible as we depended on smaller models to understand their performances. Therefore, computational resources pose a significant constraint in this work, and we want to emphasize how this factor undermines the possibility of replicating the experiments.
Here, we highlight potential threats to the validity of the research process, focusing on both external and internal factors.
External validity
When synthesizing the SynthCode2Code2NL code, we rely on code translation; we understand that synthesized data adheres to stylistic writing patterns distinct from those of humans. We tested the model’s performance on standard benchmarks. However, the impact of utilizing code snippets as synthetic data in training large language models for generalization over human text-to-code and code-to-code search is still not fully understood.
Internal validity
The ablation study focused on fine-tuning the model with and without multi-layer loss. However, this comparison does not account for how the model behaves when starting from a model not pre-trained on multi-layer loss. Although our experiments present promising results, further inspection is necessary to better understand this phenomenon.
Appendix A Synthetic dataset
SynthCode2Code2NL is a fine-tuning dataset designed for text-to-code and code-to-code search, built by augmenting CodeSearchNet Husain et al. (2019) with transpiled code snippets across multiple languages (Python, Java, Go, PHP, Ruby, C++, C, JavaScript). The dataset underwent a preprocessing phase, including deduplication based on the original and synthesized code columns. Near-deduplication was performed using Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 over character-level 5-grams to remove semantically identical snippets differing only in identifiers or function arguments.
For code-to-code search, we translated each snippet into a randomly sampled target language using the Qwen2.5-Coder-7B-Instruct model with greedy search to ensure consistency. Each dataset entry consists of a natural language description and two code snippets in different languages. SynthCode2Code2NL contains 1,071,367 samples, with original code from CodeSearchNet (Python, Java, PHP, Go) and translated code (Go, Ruby, JavaScript, Python, C++, PHP, C, Java). In Figure 5, In Figure 6 and Figure 7 some examples of code translation are shown.
Figure 5: Example of a code snippet translation from Go to Ruby.
Figure 6: Example of a code snippet translation from Java to C.
Figure 7: Example of a code snippet translation from Python to PHP.