### v2

# MoSE: Hierarchical Self-Distillation Enhances Early Layer Embeddings

Abstract
Deploying language models often requires navigating accuracy vs. performance trade-offs to meet latency constraints while preserving utility. Traditional model distillation reduces size but incurs substantial costs through training separate models. We introduce ModularStarEncoder (MoSE), a 1-billion-parameter multi-exit encoder for code retrieval and classification that employs a novel Self-Distillation mechanism. This approach significantly enhances lower-layer representations, enabling flexible deployment of different model portions with favorable performance trade-offs. Our architecture improves text-to-code and code-to-code search by targeting specific encoder layers as exit heads, where higher layers guide earlier ones during training—improving intermediate representations at minimal additional cost. We further enhance MoSE with a repository-level contextual loss that maximizes training context window utilization. Additionally, we release a new dataset created through code translation that extends text-to-code benchmarks with cross-language code-to-code pairs. Evaluations demonstrate the effectiveness of Self-Distillation as a principled approach to trading inference cost for accuracy across various code understanding tasks.
1 Introduction
Large language models (LLMs) have significantly impacted natural language processing (Niu et al., 2023), but their computational demands pose substantial deployment challenges. The field has responded with various efficiency strategies: quantization techniques reduce numerical precision (Jacob et al., 2017; Lin et al., 2023; Egiazarian et al., 2024); knowledge distillation trains smaller “student” models to emulate larger “teacher” models (Sanh et al., 2019; Jiao et al., 2019); and pruning methods eliminate less influential weights (Han et al., 2015). Concurrently, model families like LLaMA (Dubey et al., 2024), Qwen (Hui et al., 2024), Mistral (Jiang et al., 2023), and SmolLM (Allal et al., 2025) exemplify a shift toward more efficient architectures at varying parameter scales.
Dynamic inference approaches further optimize efficiency through mechanisms like multi-exit networks. Early-exit architectures such as BranchyNet (Teerapittayanon et al., 2017) balance computation and accuracy by allowing predictions at intermediate layers, while Matryoshka representation learning (Kusupati et al., 2022) enables adjustment of computational complexity through embedding dimensionality pruning. These approaches maintain performance while reducing computational demands in resource-constrained environments. Furthermore, insights from the SimCLR framework (Chen et al., 2020) revealed that the final layer of a model does not necessarily produce the most semantically meaningful representations. Valeriani et al. (2023) analyzed the geometric evolution of hidden representations within large transformer models across layers, outlining that intermediate layers are identified as holding the most semantically rich representations for downstream tasks.
[…]
Figure 1: (a) Overview of our multi-exit self-distillation encoder, shown here with exit heads at selected layers (e.g., Layers 4, 9, 18, 27, and 36). Each exit head predicts an output embedding and adds a layer loss, contribution weighted by a coefficient $\alpha_{i}$, summed into the overall objective $\mathcal{L}$. (b) Computational cost (GFLOPs) vs. performances trade-off at different exit layers over CodeSearchNet dataset (text-to-code in avg. MRR) and POJ104 (code-to-code in mAP). Despite a reduction of approximately 90% floating point operations from layer 36 to layer 4, MRR performance only drops by 6.4% in absolute terms. For POJ104, our best results are observed in the initial layers.
Building on these principles, we introduce ModularStarEncoder (MoSE), a modular multi-exit encoder built on top of the StarCoder-2 (Lozhkov et al., 2024) architecture, with 1 billion parameters that incorporate intra-model self-distillation. As illustrated in Figure 1(a), our architecture applies training objectives at multiple exit points throughout the network (layers 4, 9, 18, 27, and 36), with each exit contributing a weighted loss component to the overall objective. This approach encourages lower layers to learn better representations by mimicking higher-layer outputs. Figure 1(b) demonstrates the resulting accuracy-computation trade-off of encoders trained under this framework: despite a 90% reduction in floating point operations from layer 36 to layer 4, performance over text-to-code retrieval drops by only 6.4% in absolute terms, while the model maintains the best performances for code-to-code retrieval in very early layers.
Notably, the exit point with the best performance isn’t necessarily the final layer, as noted in previous findings due to Chen et al. (2020); Valeriani et al. (2023). This offers both deployment flexibility and a natural model selection mechanism for specific tasks. Users can select which portion of the model to use during inference, starting from the first exit point at 160M parameters. By allowing users to select an earlier exit point, MoSE provides a direct mechanism to decrease computational demands. This flexibility is crucial as smaller models generally require less energy and memory for inference, leading to more sustainable deployment options (Samsi et al., 2023).
We fine-tuned MoSE on SynthCoNL (cf. Sec. 5 for details) for text-to-code and code-to-code retrieval, and on BigCloneBench (Svajlenko & Roy, 2015) for code clone detection, achieving state-of-the-art results among open models while maintaining deployment modularity.
Our contributions are as follows:
  1. 1.
    A self-distillation methodology that trains multiple model resolutions within a unified layer stack, reducing redundancy and improving scalability—an approach that could significantly impact LLM training pipelines dependent on multiple model distillations.
  2. 2.
    MoSE, pre-trained [^1] https://huggingface.co/modularStarEncoder/ModularStarEncoder and fine-tuned [^2] https://huggingface.co/modularStarEncoder/ModularStarEncoder-finetuned, with up to 1 billion parameters and five exit points, allowing users to select model size based on memory and computational constraints.
  3. 3.
    SynthCoNL [^3] https://huggingface.co/datasets/modularStarEncoder/SynthCoNL, a new dataset of 1,071,367 natural language-code-code triplets constructed via code translation, expanding text-to-code benchmarks with cross-language code-to-code pairs.
Figure 2: The illustration on the left depicts the in-context loss framework, where samples from various repositories are concatenated. Positive examples originate from the same repository context, whereas negative examples are sourced from different repositories. To enable the model’s use of FlashAttention V2, we applied left padding and positioned the CLS token at the end of the sentence. On the right side, you’ll find the pseudocode for the in-context loss framework.
2 Self-Distillation and In-Context Classification
As different layers of transformer models capture varying degrees of semantic richness, often with intermediate layers proving most effective for specific tasks (Bolya et al., 2025), we design our pre-training strategy to support modularity and ease downstream task adaptability. Rather than treating the deepest layer as the sole source of meaningful representations, we aim to enhance representations throughout the network.
We hypothesize that applying supervision at multiple depths, not just at the output—encourages more robust and semantically rich intermediate representations. To this end, we combine multiple training objectives, each contributing to more versatile and generalizable embeddings (Wang et al., 2023). While we retain the widely adopted Masked Language Modeling (MLM) loss Feng et al. (2020), we discard the traditional Next Sentence Prediction (NSP) objective. NSP has been shown to offer minimal benefit after fine-tuning (Liu et al., 2019), and it imposes a rigid sentence-level structure that leads to inefficient use of context windows, especially for long-form code inputs. In its place, we introduce an In-Context Classification (ICC) loss (see section 2.1) that increases input density by reducing padding and fosters better chunk-level comprehension, essential for tasks like semantic retrieval or classification.
By integrating these two complementary objectives, MLM and ICC, and distributing their supervision across multiple layers, forming the basis of our Self-Distillation mechanism (see section 2.2), our training approach improves internal representation capabilities and flexibility, ultimately yielding a model better suited for modular deployment and early exiting.
2.1 Masked Language Modeling and In-Context Classification
We revisited the NSP loss and introduced an In-Context Classification (ICC) objective. We hypothesize that predicting whether multiple code snippets belong to the same context (in our case, the same repository) can enhance semantic search performance while allowing efficient concatenation of multiple code fragments. Our final training objective is the summation of two losses: (1) MLM loss and (2) ICC loss: $\mathcal{L}=\mathcal{L}_{MLM}+\mathcal{L}_{ICC}$.
[…]
Table 1: Performance—reported in MRR, and when specified in NDCG—of different models on text-to-code with CodeSearchNet using CodeXGLUE. We reported the results presented in CodeT5+, UniXcoder and ModernBERT (Wang et al., 2023; Guo et al., 2022; Warner et al., 2024). MoSE surpasses all existing open-source models, achieving an improvement of +3.6% compared to the best-performing open model, CodeT5+. MoSE reduced the performance gap between open-source models and closed-source ones such as OpenAI text-embedding-3-large.
CodeSearchNet
Model Ruby JS Go Python Java PHP avg. MRR avg. NDCG
MoSE 74.1 74.0 82.5 92.5 78.7 84.5 81.0 84.2
CodeT5+ (770M) 78.0 71.3 92.7 75.8 76.2 70.1 77.4 -
UniXcoder 74.0 68.4 91.5 72.0 72.6 67.6 74.4 -
ModernBERT-large - - - - - - - 59.5
OpenAI text-embedding-3-large 84.7 85.3 95.9 99.8 90.1 95.6 91.9 93.3
2.2 Self-Distillation: Multi-Layer loss
To achieve layer-wise modularity in transformer architectures, we apply the previously introduced loss (section 2.1) across a selected set of layers, sharing classification heads (masked language modeling and in-context classification) while incorporating a positional embedding of the layer index. The total loss is computed as the sum of individual layer losses, weighted by a factor $\alpha$ to prioritize deeper layers: $\mathcal{L}=\sum_{i\in\iota}\mathcal{L}_{i}\cdot\alpha$ where $\alpha=i/|I|$ and $I=\{1,\dots,36\}$ represents all layers, and the selected subset $\iota=\{4,9,18,27,36\}$ defines the layers where the loss is applied. The selected subset was chosen to enable four model variants equally spaced in depth (9, 18, 27, 36) along with an additional “tiny” version (4) to see the model performance in a lower number of parameters set. This Self-Distillation mechanism, shown in Figure 1, allows for flexible model deployment, enabling adaptive layer pruning with minimal impact on performance.
3 Pre-training
We built MoSE on top of StarCoder-2 (Lozhkov et al., 2024), applying several modifications to the model. We reduced its original size, obtaining a model of 1B parameters. Our architecture comprises 36 hidden layers and adopts Grouped Query Attention (GQA) (Ainslie et al., 2023) with 16 attention heads and 4 key-value heads. MoSE relies upon Rotary Positional Encoding (RoPE) (Su et al., 2021) with a base period $\theta=10^{-6}$ and features a hidden dimensionality of 1,024 with an intermediate size of 12,288.
To make our model class bidirectional, we removed the causal masking from the self-attention operations within StarCoder-2. Aiming for modularity, we also replaced sliding window attention with full attention. This step was taken to avoid the receptive field phenomenon of the sliding window mechanisms (Zhu et al., 2021). Finally, our implementation integrates FlashAttention V2 (Dao, 2023) for faster inference. We have included and summarized additional details about model architecture in appendix D.
3.1 Training Details
We pre-trained MoSE with a batch size of 4M tokens with a maximum context length of 2,048 tokens, for 245,000 training steps, processing approximately 1T tokens from TheStackV2 (Lozhkov et al., 2024) dataset.
We used the AdamW optimizer with $\beta_{1}$ set to 0.9, $\beta_{2}$ to 0.95, $\epsilon$ to $1\mathrm{e}{-6}$, and a weight decay of $1\mathrm{e}{-1}$. We initialized the learning rate at $6.24\mathrm{e}{-4}$ and decreased it using a multi-step learning rate scheduler (Bi et al., 2024) with 4,000 warmup steps. The learning rate was reduced at 120,000, 185,000, 220,000, 230,000, and 240,000 training steps, applying a decay factor of 0.36, and from step 185,000 onward, further reduced by factors of 0.1, 0.031, 0.01, and 0.001. We conducted pre-training and fine-tuning on 512 NVIDIA Ampere (64GB) GPUs, requiring 450,000 GPU working hours on the Leonardo supercomputer (Turisini et al., 2023).
4 Downstream tasks
4.1 Finetuning for Retrieval, text-to-code and code-to-code search
Using the SynthCoNL dataset detailed in section 5, we leveraged a instruction prompting approach Su et al. (2023) and trained a single model for both text-to-code and code-to-code retrieval tasks. The optimization objective uses CLIP loss (Radford et al., 2021) with the same multi-layer approach discussed in section 2.2, modified to enhance representation learning. Accordingly, we replaced the single-head projection of the multi-layer loss with five distinct projection heads, applied at different exit points of the pre-trained model (layers 4, 9, 18, 27, and 36). We used a batch of 2,048 elements, ensuring that text-to-code and code-to-code were equally distributed across the batch.
We performed data augmentation by randomly replacing frequently occurring words (appearing more than twice and having at least three characters) with random strings. We applied the augmentation exclusively to code snippets in 30% of cases, leaving natural language descriptions unchanged. After conducting a grid search, we selected $1\mathrm{e}{-5}$ as the learning rate, maintained throughout the finetuning process, and set the temperature parameter at 10.0.
Table 2: BigCloneBench code clone detection performance in a classification setup. MoSE maintains good performance (%) throughout different layers (L) with results on par with the state-of-the-art.
BigCloneBench
Model Recall Precision F1
MoSE (L4) 96.4 89.8 93.0
MoSE (L9) 96.6 90.4 93.4
MoSE (L18) 96.4 92.1 94.2
MoSE (L27) 96.5 91.8 94.1
MoSE (L36) 96.5 91.7 94.1
UniXcoder 92.9 97.6 95.2
CodeBERT 94.7 93.4 94.1
CodeT5+ (770M) 96.7 93.5 95.1
4.2 Finetuning for classification, code clone detection
[…]
5 SynthCoNL
Prompt Translate this “‘print("Hello World")”’ from Python to Rust. Here is the translated code: Qwen2.5-Coder-7B-Instruct Output fn main() { println!("Hello World!"); }
Figure 3: Prompt provided to Qwen2.5-Coder-7B-Instruct for translating a given code snippet ( print("Hello World") in the example) from a source programming language (Python) to a target one (Rust).
We built SynthCoNL, a dataset that supports training and evaluation of encoders in text-to-code and code-to-code search, under multiple languages. Using the popular CodeSearchNet (Husain et al., 2019) as a seed dataset and selecting popular programming languages (Python, Java, Go, and PHP), we augmented it by transpiling available code snippets onto other languages.
To generate semantically similar code snippets for code-to-code search, we translated each snippet into a different language randomly sampled from Go, Ruby, Python, Java, C++, PHP, C, JavaScript. We prompted the Qwen2.5-Coder-7B-Instruct model with the source code, the name of the source language, and the name of the target language (see fig. 3). During code translation, we carried out a greedy sampling so as to prevent semantic discrepancies.
This process yielded pairs of code snippets in distinct languages tied to the same natural language description. As a result, every sample in the fine-tuning dataset includes a natural language description and two code snippets from distinct languages. SynthCoNL contains 1,071,367 triplets where, in the first code column, we directly sampled code snippets from CodeSearchNet, including Python, Java, PHP, and Go. The second code column, artificially synthesized via code translation, includes Go, Ruby, JavaScript, Python, C++, PHP, C, and Java code snippets. After a manual inspection, we discovered that both columns contained code snippets that differed only in identifiers or function arguments. Several tasks were semantically identical but paraphrased with different parameter requirements (e.g., two identical paraphrased tasks asked for opening a socket on a different port). During the preprocessing phase of SynthCoNL, motivated by the dataset’s redundancy and preliminary experiments that show its effectiveness on the model’s performance, we near-deduplicated the dataset using both the CodeSearchNet code column and the synthesized code column. During the data near deduplication phase, we relied on Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 and 256 permutations, analyzing character-level 5-grams.
6 Results and Discussion
6.1 Evaluation
Figure 4: Performance of different models on CT and POJ104 for code-to-code retrieval with CodeXGLUE dataset. MoSE has state-of-the-art results with the CT benchmark and outperforms all the open-source models over the POJ104 benchmark. Compared to closed-source solutions (OpenAItext-embedding-3-large), MoSE performs on par in the code translation task and strongly reduces the gap between open- and closed-source models with the POJ104 benchmark.
We evaluated MoSE, after fine-tuned for retrieval, on diverse code understanding tasks, leveraging the CodeXGLUE (Lu et al., 2021) group of benchmarks.
On CodeSearchNet dataset, we assessed cross-modal (natural language to code) learning by retrieving target code from 999 distractors using natural language queries. For code-to-code retrieval, again against 999 distractors, we utilized CodeXGLUE’s Code Translation (CT) dataset to test cross-lingual retrieval (e.g., Java to C# snippets implementing the same functionality), and its POJ104 dataset (C++ snippets, semantically equivalent but syntactically different) for intra-lingual semantic search, evaluating generalization over structural differences while preserving semantics. In our retrieval performance evaluation, when comparing MoSE with other encoders, we included results from OpenAI’s text-embedding-3-large as a representative commercial, closed-source model.
Additionally, MoSE was fine-tuned and evaluated on the separate BigCloneBench (Svajlenko & Roy, 2015) dataset for Java code clone detection (a binary classification task), assessing its classification performance.
6.2 Benchmarks
Table 1 presents the results for the CodeSearchNet (text-to-code) task in terms of Mean Reciprocal Rank (MRR, higher is better) for each single language, average Normalized Discounted Cumulative Gain (NDCG, higher is better), and average MRR. Results for UniXcoder, ModernBERT, and CodeT5+ are reported from the original papers (Guo et al., 2022; Warner et al., 2024; Wang et al., 2023). On CodeSearchNet, MoSE achieves an MRR of 81.0 and a NDCG of 84.2, outperforming CodeT5+ (770M), UniXcoder, and ModernBERT-large. The only encoder that surpasses MoSE is OpenAI’s text-embedding-3-large.
In Figure 4, we present results from both POJ104 and CT datasets reported respectively in terms of MRR for code translation (Java to C# retrieval) and mean average precision for POJ104 (C++ to C++ retrieval). MoSE reaches the best performance among the evaluated models. We further replicated the benchmarking for all models in a zero-shot setting for code-to-code tasks as our model does not integrate POJ104 and the code translation datasets in the training set. We present a comprehensive and detailed analysis of code retrieval errors in appendix C.
Figure 5: Similarity score Heatmap presenting results from permutation tests (10,000 permutations with $\alpha$ = 0.05) across different exit points for Python-to-Java retrieval. * indicates p < 0.05 and ** indicates p < 0.001. Different layers, despite a common training objective, yield different similarity scores.
Referring to Figure 4, on the POJ104 dataset in zero-shot, MoSE achieves a mean Average Precision (mAP, higher is better) of 75.9, which is state-of-the-art among open-sourced models. However, it is significantly behind OpenAI text-embedding-3-large. We underscore that a direct comparison with OpenAI text-embedding-3-large remains challenging because it is closed-source, and details such as model size, training methodology, or potential data contamination are undisclosed.
We observe a marked difference in performance depending on whether the task is framed as text-to-code or code-to-code retrieval. As shown in Figure 1(b), shallow layers yield better performance in the code-to-code setting, whereas for text-to-code retrieval, optimal performance emerges at layer 18.
This representation divergence suggests that the input modality influences where semantic alignment is most effectively captured within the model. This observation aligns with Valeriani et al. (2023) findings, underlining that semantic information is distributed unevenly across layers and is better expressed in the intermediate ones. MoSE leverages this property by supporting dynamic exit points, enabling users to balance computational cost and task performance based on where the most relevant representations emerge.
To further investigate this phenomenon, we compared similarity scores across layers to assess how each layer represents code snippets differently. Figure 5 shows permutation test results for a Python-to-Java retrieval task, highlighting significant variations in similarity scores that support our previous findings. A detailed analysis of these results is provided in appendix C.
In Table 2, we report classification performance across different exit layers of MoSE for the BigCloneBench benchmark. Results show stable performance throughout layers, with F1-scores from Layer 18 onward (94.2–94.1) closely matching those of deeper exits. Earlier layers (e.g., Layers 4 and 9) exhibit slightly lower precision, reflecting the expected trade-off between early inference and representational depth.
Our method remains competitive compared to state-of-the-art models. While CodeT5+ (770M) and UniXcoder achieve peaks F1 of 95.1 and 95.2, MoSE delivers on-par results (-0.9) by Layer 18, with the added benefit of a modular architecture.
6.3 Ablation Study
We conducted an ablation study for the Multi-Layer loss, by fine-tuning (starting from MoSE, pre-trained) on SynthCoNL, each exit point singularly (e.g., for the baseline on layer 18, we retain only the first 18 layers and fine-tune the model using just the projection head on that layer). Finally, we compared the baseline models with the corresponding results (Self-Distilled) of MoSE fine-tuned for retrieval with the multi-layer loss; all tests are conducted on the CodeSearchNet dataset. MoSE consistently outperforms the single-exit baselines (up to +4.36% averaged Recall@1 gain on layer 9), indicating that lower-level layers benefit from training signals propagated from deeper layers. This behavior is highlighted in Table 3, where MoSE, indicated as Self-Distilled, outperforms all the single exit baselines consistently. This finding underscores a promising new direction in Self-Distillation for large-scale code and text models, enabling high performance even in more compact configurations. Moreover, Figure 6 illustrates that MoSE maintains robust performance from layers 18 to 36, allowing users to scale down the network to match their memory, computational, or latency constraints while preserving strong retrieval performances.
Table 3: Performance comparison of MoSE layers and baseline fine-tuned models on the CodeSearchNet benchmark. The table displays the overall retrieval performance measured by the Mean Reciprocal Rank (MRR, higher is better). We refer to MoSE, fine-tuned with multiple exit points simultaneously, as Self-Distilled. The models not marked as self-distilled are the baselines, fine-tuned individually for each exit point. MoSE fine-tuned with multi-layer loss (Self-Distillation) consistently outperforms the baselines, showing a more significant difference in the initial layers.
CodeSearchNet
Model type Num. Layers Num. Params. Ruby Javascript Go Python Java PHP avg. MRR
Base 4 $160M$ 59.5 61.3 72.1 86.2 68.2 75.5 70.5
Self-Distilled 62.2 64.7 74.8 88.1 71.4 78.0 73.2
Base 9 $300M$ 64.9 65.7 74.3 87.3 72.0 78.8 73.8
Self-Distilled 67.6 69.4 78.9 90.2 75.5 82.3 77.3
Base 18 $550M$ 73.8 73.5 82.4 92.1 78.4 84.0 80.7
Self-Distilled 74.1 74.0 82.5 92.5 78.7 84.5 81.0
Base 27 $800M$ 72.3 71.8 80.8 90.8 76.9 82.3 79.1
Self-Distilled 73.2 73.3 81.7 92.1 77.8 83.8 80.3
Base 36 $1B$ 72.3 72.9 80.7 91.5 77.1 82.9 79.5
Self-Distilled 73.5 72.6 80.5 91.4 76.9 82.7 79.6
7 Related work
Figure 6: Performance Comparison between MoSE fine-tuned with Multi-layer loss (Self-Distilled model) and baselines fine-tuned just on single exit points: The graph illustrates averaged MRR and Recall@1 results for different layers over the CodeSearchNet test set. Layers fine-tuned with Multi-Layer loss outperforms the baselines (up to 3.47% MRR and +4.36% Recall@1).
Since the introduction of ELMo (Peters et al., 2018), deep contextual information has enhanced generating embeddings for textual retrieval or classification, reaching state-of-the-art results in several tasks. BERT (Feng et al., 2020) followed those findings, adapting the Transformer architecture (Vaswani et al., 2017) to enable a bi-directional representation with two different training objecting, namely the masked language modeling and the next sentence prediction losses. Lan et al. (2019) and Liu et al. (2019) adapted the BERT architecture to obtain an enhanced pre-trained model by removing or modifying the NSP, focusing on pre-training data or hyperparameters optimization. More recently, modernBERT (Warner et al., 2024) tied the gap between modern decoders (Jiang et al., 2023; Hui et al., 2024; Dubey et al., 2024; Touvron et al., 2023; Lozhkov et al., 2024) advancements that rely upon models with an increased number of parameters, trained upon more tokens, and being capable of handling longer contextual information.
In code representation, large language models must be adapted by training them on a curated corpus focused on software and by leveraging code’s syntactic and semantic structures, which differ significantly from natural language. Feng et al. (2020) adapted the BERT architecture to produce semantically meaningful embeddings for source code, resulting in codeBERT. This was accomplished by including more source code in the training set and focusing on a training loss that can leverage bimodal (natural language and code) contextual information (Clark et al., 2020). GraphCodeBERT enhanced codeBERT (Feng et al., 2020) representations by incorporating data flow graphs, capturing dependencies between variables and operations, and improving tasks like code summarization and clone detection. UniXcoder (Guo et al., 2022) extended this by introducing a unified encoder-decoder framework, integrating abstract syntax trees (ASTs) and data flow information. Wang et al. (2023) expanded these findings with CodeT5+, stressing how multiple losses that leverage code semantics impact the model pertaining. The work incorporated text-code contrastive learning, text-code Matching, and text-code causal LM for better code understanding and generation.
When trying to achieve better performance, research has shifted toward models with a high number of parameters. While this trend appears effective from a performance perspective, end users may face computational or memory limitations as LLMs vary from millions to billions of parameters. Sanh et al. (2019) pioneered the introduction of knowledge distillation, using a “teacher” model that guides a smaller model to emulate its behavior. This methodology has been widely adopted and improved upon recently (DeepSeek-AI et al., 2025; Hui et al., 2024), becoming a standard for obtaining high-performing smaller LLMs.
Our work differs from previous work by adapting a modern architecture (StarCoder-2 (Lozhkov et al., 2024)) to a code encoder-only based model and introducing a novel ’Self-Distillation’ mechanism. We replace the next sentence prediction loss with an In-Context Classification focused on the repository level and expand the context to 2,048 tokens. Our novel Self-Distillation mechanism improves low-level layers, resulting in a modular transformer architecture without additional teacher models or further data for distillation.
8 Conclusion
We introduced MoSE, a modular, multi-exit encoder that uses Self-Distillation to enhance early-layer representations while achieving state-of-the-art results on code understanding benchmarks. By integrating multiple exit points, MoSE enables explicit efficiency-accuracy trade-offs tailored to deployment constraints. Our approach not only allows for modular use but also enhances the alignment between task requirements and representation depth. This improves task-specific layer specialization, providing the end-user with a more adaptable model.
Future work will explore combining exit-points and further understanding task type and depth interactions. While our results are encouraging, we acknowledge that computational constraints limited extensive hyperparameter tuning and pre-training ablations, and that fine-tuning with synthetic code generated from code translation has an unclear impact. We discuss these limitations in detail in appendix A.
Acknowledgments
We acknowledge ISCRA for awarding this project access to the LEONARDO supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CINECA (Italy).
This work was supported by Future AI Research (FAIR) PE01, SPOKE 8 on PERVASIVE AI funded by the National Recovery and Resilience Plan (NRRP).
Appendix A Limitations
Due to our dependence on multiple GPUs, we encountered significant computational constraints. Parameter grid searches using smaller and earlier models were the only methods to extrapolate the best hyperparameter setup. The optimal hyperparameters for smaller models can differ from those for larger ones; consequently, we encountered a limitation in discovering an effective training setup. Conducting an ablation study for the in-context classification loss was not practicable due to the high costs of the pre-training stage. Therefore, computational resources pose a significant constraint in this work, and we want to emphasize how this factor can also undermine the possibility of replicating the experiments.
Here, we highlight potential threats to the validity of the research process, focusing on both external and internal factors.
External validity
When synthesizing the SynthCoNL code, we rely on code translation; we understand that synthesized data adheres to stylistic writing patterns distinct from those of humans. We tested the model’s performance on standard benchmarks. However, the impact of utilizing code snippets as synthetic data in training large language models for generalization over human text-to-code and code-to-code search is still not fully understood.
Internal validity
The ablation study focused on fine-tuning the model with and without multi-layer loss. However, this comparison does not account for how the model behaves when starting from a model not pre-trained on multi-layer loss. Although our experiments present promising results, further inspection is necessary to better understand this phenomenon.
Appendix B Synthetic dataset
Table 4: SynthCoNL composition: Average character count and sample size per language for the CodeSearchNet seed dataset and synthesized portion obtained through translation.
Language CSN samples CSN avg. char Synth. samples Synth. avg. char
English 1,071,367 180 - -
PHP 280,706 514 116,967 579
Python 274,454 474 117,374 518
Go 234,089 350 124,125 541
Java 282,118 505 116,098 707
C++ - - 141,956 938
Ruby - - 158,494 456
C - - 136,365 1,029
JavaScript - - 159,988 557
SynthCoNL is a fine-tuning dataset designed for text-to-code and code-to-code search, built by augmenting CodeSearchNet Husain et al. (2019) with transpiled code snippets across multiple languages (Python, Java, Go, PHP, Ruby, C++, C, JavaScript). The dataset underwent a preprocessing phase, including deduplication based on the original and synthesized code columns. Near-deduplication was performed using Locality Sensitive Hashing (LSH) with a Jaccard similarity threshold of 0.7 over character-level 5-grams to remove semantically identical snippets differing only in identifiers or function arguments.
For code-to-code search, we translated each snippet into a randomly sampled target language using the Qwen2.5-Coder-7B-Instruct model with greedy search to ensure consistency. Each dataset entry consists of a natural language description and two code snippets in different languages. SynthCoNL contains 1,071,367 samples, with original code from CodeSearchNet (Python, Java, PHP, Go) and translated code (Go, Ruby, JavaScript, Python, C++, PHP, C, Java). In Figure 11, In Figure 12 and Figure 10 some examples of code translation are shown.
Table 4 shows the average number of characters per language in SynthCoNL, we emphasize that synthesized data is significantly longer than human-written code and might have stylistic differences compared to human code, we further discussed this in the Limitations section.
Appendix C Failure Case Analysis in Code-to-Code Retrieval
Figure 7: Retrieved incorrect samples from CodeT5+, UniXcoder, and the 18th layer of MoSE. All the models appear to be misled by structural similarities, such as loop functions and return values, highlighting a common superficial retrieval phase.
Figure 8: Histograms of normalized similarity score between correct and incorrect retrieved code snippets. The first 5 distributions belong respectively to MoSE 4, 9, 18, 27, 36 layers’ exit points; CodeT5+ and UniXcoder are the last ones displayed. While our model tends to have different, distinct distributions regarding wrong and correct retrievals, both UniXcoder and CodeT5+ have more overlapping results.
Table 5: Permutation tests between correct and incorrect similarity scores ($\alpha=0.05$). The table shows a big discrepancy between MoSE t-statistics and both UniXcoder and CodeT5+ results.
Model t-statistic p-value
MoSE (L4) 8.0 <0.001
MoSE (L9) 7.3 <0.001
MoSE (L18) 8.0 <0.001
MoSE (L27) 7.8 <0.001
MoSE (L36) 6.3 <0.001
UniXcoder 2.2 0.04
CodeT5+ 110M-embedding 2.7 0.009
Figure 9: Heatmap of permutation tests (10,000 permutations) between MoSE similarity score results ($\alpha=0.05$). * with p<0.05 and ** with p<0.001. In the left heatmap, we present the differences in correctly retrieved snippets across various layers, while the right heatmap illustrates the incorrectly retrieved snippets. The heatmaps demonstrate that even layers trained to achieve similar representation objectives can yield different similarity scores.
In this section, we analyzed where our model fails in code-to-code retrieval, relying upon gfg, a Python-to-Java dataset. We compared how our model gets misled by different code snippets with CodeT5+ and UniXcoder, and the confidence level with which the model identifies the occurrences. In fig. 7 we highlight a common defect of the most used open code encoders. When retrieving code snippets focusing on semantic similarity, we emphasize how all the inspected models might get misled by structural factors such as loops, if statements, and return types. The retrieved results underline an overconsideration of structural features despite semantic differences.
We then analyzed whether the model’s confidence might suggest uncertainty in retrieving incorrect snippets of code, indicating a statistically significant level of "awareness" in overconsidering misleading features. Thus, we compared correct and incorrect similarity distribution scores for retrieved code snippets. Figure 8 shows how MoSE, CodeT5+, and UniXcoder results differ when comparing correct and incorrect retrievals. In table 5, MoSE demonstrated a more significant t-statistic than CodeT5+ and UniXcoder. In a retrieval use case scenario, when the confidence level is insufficient, we suggest that thresholding the similarity score could be an option when deciding whether to provide the end user with a single result or a herd of results.
In fig. 9, we present the differences in similarity scores among the five exit points using a permutation test. It is important to note that, although the model is trained to maintain similar retrieval capabilities across these five exit points (as discussed in section 2.1), and while the incorrectly retrieved snippets might overlap between the layers, each exit point encodes code snippets differently.
Based on these results, we strongly recommend conducting a grid search between the similarity score threshold and the exit point to utilize.
Figure 10: Example of a code snippet translation from Python to PHP.
Figure 11: Example of a code snippet translation from Go to Ruby.
Figure 12: Example of a code snippet translation from Java to C.
Appendix D Model card
Table 6 provides a comprehensive overview of the hyperparameters, architectural choices, and training configurations used in the development of MoSE. This includes detailed settings for both the pre-training and fine-tuning stages, as well as hardware specifications. The information is provided to enhance reproducibility and improve understanding of the model’s design.
Table 6: Details on hyperparameters, architecture, and training for MoSE are presented here, including insights from both pre-training and fine-tuning stages.
Architecture Pre-training
Hyperparameter Value Hyperparameter Value
Model size 1B parameters Batch size 3.99M tokens
Precision bfloat16 Pretraining steps 245,000
Hidden layers 36 Pretraining tokens 1T
Attention heads 16 Loss function MLM + In-Context loss
Hidden dimensionality 1,024 Multi-layer loss Yes
Positional encoding RoPE ($\theta=10^{-6}$) Optimizer AdamW
Context length 2,048 Weight decay $1\mathrm{e}{-1}$
Attention mechanism Grouped Query Attention Initial learning rate $6.24\mathrm{e}{-4}$
Attention pattern Bi-directional Learning rate schedule Multi-step
Warmup steps 4,000
Fine-tuning Hardware (Pre-training + Fine-tuning)
Hyperparameter Value Hyperparameter Value
Dataset size 635,404 samples Pre-training GPUs 512 NVIDIA Ampere A100 (64GB)
Fine-tuning steps 20,000 Fine-tuning GPUs 512 NVIDIA Ampere A100 (64GB)
Loss function CLIP loss Benchmarking GPUs 1 NVIDIA Ampere A100 (64GB)
Multi-layer loss Yes Overall training hours 450,000
Batch size 2,048
Learning rate $1.0\mathrm{e}{-5}$
Temperature parameter 10.0